Sketch the region enclosed by the curves and find its area.
9. J = ਸੰਢ, 3 = z| 10. y = 3 = 2
Answer #1


\small {\color{Red} (9)}\\ \text{Consider the curves}\\ y=\frac{2}{1+x^2} \: \: \: y=\left | x \right |\\ \text{The absolute function is as follows}\\ y=\left | x \right |=\begin{cases} x & \text{ if } x\geq 0 \\ -x & \text{ if } x < 0 \end{cases}\\ \text{The point of intersection of the two curves is as follows}\\ \frac{2}{1+x^2}=x\Rightarrow x=1\\ \frac{2}{1+x^2}=-x\Rightarrow x=-1\\ \text{Therefore, the point of intersection are (-1,1) and (1,1)} \\

\small \text{The area between the two curves is as follows}\\ \mathrm{Area}=\int_{a}^{b}\left [ f(x)-g(x) \right ]dx\\ =\int_{-1}^{0}\left [ \frac{2}{1+x^2}-(-x) \right ]dx+\int_{0}^{1}\left [ \frac{2}{1+x^2}-(x) \right ]dx\\ =\int_{-1}^{0}\left [ \frac{2}{1+x^2}+x \right ]dx+\int_{0}^{1}\left [ \frac{2}{1+x^2}-x \right ]dx\\ =\left[2\arctan \left(x\right)+\frac{x^2}{2}\right]^0_{-1}+\left[2\arctan \left(x\right)-\frac{x^2}{2}\right]^1_{0}\\ =\left ( \frac{\pi }{2}-\frac{1}{2} \right )+\left ( \frac{\pi }{2}-\frac{1}{2} \right )\\ =2\left ( \frac{\pi }{2}-\frac{1}{2} \right )\\ =\pi-1\\ =2.14\\

\small {\color{Red} (10)}\\ \text{Consider the curves}\\ y=\frac{1}{\sqrt{1-x^2}} \: \: \: y=2\\ \text{The point of intersection of the two curves is as follows}\\ \frac{1}{\sqrt{1-x^2}}=2\\ \sqrt{1-x^2}=\frac{1}{2}\\ 1-x^2=\frac{1}{4} \Rightarrow x=\frac{\sqrt{3}}{2},\:x=-\frac{\sqrt{3}}{2}\\ \text{Therefore, the point of intersection are }\left (-\frac{\sqrt{3}}{2},2 \right ) ,\; \left (\frac{\sqrt{3}}{2},2 \right )\\ \text{The area between the two curves is as follows}\\ \mathrm{Area}=\int_{a}^{b}\left [ f(x)-g(x) \right ]dx\\ =\int_{-\frac{\sqrt{3}}{2}}^{\frac{\sqrt{3}}{2}}(2-\frac{1}{\sqrt{1-x^2}})dx\\ =\left[2x-\arcsin \left(x\right)\right]^{\frac{\sqrt{3}}{2}}_{-\frac{\sqrt{3}}{2}}\\ =2\sqrt{3}-2\cdot \frac{\pi }{3}\\ =1.3697\\